OpenStudy (anonymous):

4(x - 6) = -2(x - 3) Solve?

OpenStudy (anonymous):

Distribute: 4x-24=2x-6 Rearrange so that the "x" is on one side of the equal sign and rest on the other: 4x-2x=24-6 (keep in mind that when you move them over, the sign changes) Basic adding/subtration: 2x=18 I believe you can solve the rest now.

OpenStudy (anonymous):

Whoa! Thanks a bunch! :D